OpenStudy (lord2clash):

Which set of three numbers could be the side lengths of a triangle? A. 4, 10, 6 B. 4, 7, 7 C. 4, 10, 3 D. 4, 7, 11

OpenStudy (lord2clash):

@pooja195 @Love.Bug

pooja195 (pooja195):

If a+b is greater than c you can make a triangle if its less you cant make a triangle

OpenStudy (love.bug):

i think D but you would need a^2xb^2=c^2

OpenStudy (anonymous):

its d

OpenStudy (lord2clash):

okay thats what i was thinking but pooja said if its greater so i was confused... well thank you cx

pooja195 (pooja195):

Why don't you guys explain your answers rather than just give them away?

pooja195 (pooja195):

Explain WHY they are correct/incorrect.

OpenStudy (igreen):

Pooja meant if you add the two smaller numbers given, and it's greater than the last bigger number, you can make a triangle with them.

pooja195 (pooja195):

D is incorrect...

OpenStudy (igreen):

For example. A. 4, 10, 6 Two smaller numbers: 4, 6 Add them 4 + 6 = 10 10 is not greater than 10, therefore you cannot make a triangle with these three measurements.

OpenStudy (anonymous):

that is i good idea @pooja195 most of the time i do but you did it for me.

OpenStudy (love.bug):

haha i knew it hold on let me figure out what it is real quick lol @pooja195

OpenStudy (lord2clash):

.... D was indeed incorrect .... my apex says the correct answer was B. 4,7,7

OpenStudy (love.bug):

ahhhh i see where i went wrong...sorry!

OpenStudy (igreen):

\(\sf a^2 + b^2 = c^2\) Where \(\sf a\) and \(\sf b\) are the shorter sides and \(\sf c\) is the longer side(hypotenuse).

OpenStudy (anonymous):

It's "b" The lengths of the sides of a triangle; the sum of each pair must be greater than the remaining side.

OpenStudy (lord2clash):

well thank you guys for helping cx